OpenStudy (anonymous):

Make a conjecture about 8Tn + 1 and then prove it. Triangular numbers Tn. Hint, try several examples using numbers to get a pattern and make a conjecture

OpenStudy (anonymous):

Well you could go that 8Tn + 1 is odd for all n

OpenStudy (wolf1728):

The formula for a triangular number is n * (n+1) /2 So the 4th triangular number equals 4 * (4+1) / 2 = 20 / 2 equals 10

OpenStudy (anonymous):

To prove, clearly all Tn are integers, and so 8Tn must be even. Adding one therefore must give odd.

OpenStudy (anonymous):

Hope this is what you are looking for

OpenStudy (anonymous):

\[8T{n}+1\] I am looking for a conjecture dealing with triangular numbers

OpenStudy (anonymous):

isnt what i suggested a conjecture then?

OpenStudy (anonymous):

Yes but they are meaning another pattern that equals the same such as (2n+1)^2 is the same as 8Tn +1, but I am trying to prove it

OpenStudy (anonymous):

Ah why didnt you say so earlier

OpenStudy (anonymous):

So you want to prove (2n + 1)^2 = 8Tn + 1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You know Mathematical Induction?

OpenStudy (anonymous):

Actually dw

OpenStudy (anonymous):

Ok can we assume Tn = n(n+1)/2? Or do i have to prove that...

OpenStudy (anonymous):

Ok I'm gonna assume we can, and prove it later if i get time

OpenStudy (anonymous):

I know how to do induction. I can not figure it out at n = k+1

OpenStudy (anonymous):

So on the LHS we have (2n+1)^2, and RHS we have 8Tn + 1 work on RHS, RHS = 8*n*(n+1)/2 + 1 = 4n(n+1) + 1 = 4n^2 + 4n + 1 = (2n+1)^2 expand it if you don't believe me

OpenStudy (anonymous):

= LHS Problem Solved

OpenStudy (anonymous):

All good?

OpenStudy (anonymous):

May I get the entire induction without the LHS and RHS